OpenStudy (anonymous):

4e-8 = 32 please show how you did this problem?

OpenStudy (anonymous):

4e = 32 + 8 4e = 40 e = 10

OpenStudy (anonymous):

how did you solve the problem?

OpenStudy (anonymous):

4e-8 = 32 first u must get rid of the 8 by adding 8 to each side.then get rid of the 4 by dividing since it is 4e (4 times e) and do the same to the other side so its e= 10.

OpenStudy (anonymous):

4e - 8 = 32 4e - 8 + 8 = 32 + 8 4e = 32+8 4e = 40 4e 40 ---- = ---- 4 4 e = 10

OpenStudy (anonymous):

do u need help